# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is highly inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the cache. This enables building what we call intrinsics, i.e. highly specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We use aLoRA to train a set of intrinsics models, demonstrating competitive accuracy with standard LoRA while achieving significant inference benefits. The codebase is at https://github.com/IBM/activated-lora.
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands fine-tuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA)Hu et al. (2021) has addressed these challenges by introducing a parameter-efficient method for fine-tuning Houlsby et al. (2019), enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a lightweight and effective alternative for task-specific customization, particularly for large foundation models. Yet, while LoRA excels in static or single-task scenarios, it presents inefficiencies when applied in multiturn interactions, where dynamic switching between multiple LoRA configurations is required in agentic or reasoning pipelines.
@@
This article is a sister companion to Danilevsky et al. (2025), which focuses on the development of the RAG intrinsics for which we provide corresponding aLoRA implementations.
@@ 1.1 Background
This lowers the number of parameters, making finetuning significantly more efficient Hu et al. (2021). If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then the KV cache for the input must be recomputed.
@@ 2 Intrinsics and late prompts
Inference on LLMs is becoming quite sophisticated; for example, sending prompts directly to LLMs is a largely deprecated practice by developers, with structured interfaces (e.g. tool descriptions, and a sequence of messages from a variety of roles and with various content types) being processed by lower level, model specific software or template configurations to translate such structured interface to actual tokens. Thus we posit LLM intrinsics are best implemented by the combination of a model and a software layer that leverages advances on model inference technologies to provide differentiated performance through a stable interface.
@@
[…]
Figure 2: Late vs. early prompting framework for intrinsics. The aLoRA adapter architecture is designed to preserve the cache-reuse benefits of late prompting by adapting weights only on the blue tokens, allowing it to reuse the base model cache for the red input tokens.
@@ 3.1 Invocation
3.2 Comparison to LoRA at inference time
In Figure 3, an example is shown that conforms to the late prompting framework. As part of an ongoing base model chat, a user query comes in, and (optionally) documents or other information is also included. The base model then generates a response. We wish to use Certainty and Hallucination intrinsics to check this base model response, so we append the corresponding invocation sequences and generate the intrinsics outputs. These intrinsics calls can be done in parallel as they do not depend on each other. For the LoRA-based intrinsics, the entire context must be prefilled using the LoRA model prior to being able to generate the intrinsic output. On the other hand, Activated LoRA can reuse the base model cache for the vast majority of the input, only needing to prefill its own short prompt. This results in signficant latency gains.
Refer to caption (b)
Figure 3: LoRA (a) vs. aLoRA (b) intrinsic generation calls, checking an output generated by the base model. Note that aLoRA can reuse the long base model cache, saving signficant prefill costs in both compute and GPU memory.
3.3 Training and Inference
Our implementation of aLoRA at https://github.ibm.com/Kristjan-H-Greenewald/activated-lora/tree/main is designed to seamlessly use standard Huggingface libraries Wolf et al. (2019) for training and inference (using the KV cache).
@@ 5 Conclusion
Future work will explore integrating aLoRA into modern inference servers such as vLLM and developing expanded use cases for aLoRA models in agentic and test-time-scaling applications.
@@ B.1 QA datasets for Uncertainty Quantification Intrinsic
B.2 Training data for Hallucination Detection Intrinsic
The following public datasets were used for finetuning. The details of data creation for RAG response generation are available in the Granite Technical Report Granite Team [2024]. For creating the hallucination labels for responses, the technique described in Achintalwar et al. [2024] was used.
B.3 Training data for Query Rewrite Intrinsic
@@
B.4 Training data for Answerability Determination Intrinsic

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is highly inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the cache. This enables building what we call intrinsics, i.e. highly specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We train a set of aLoRA-based intrinsics models, demonstrating competitive accuracy with standard LoRA while achieving significant inference benefits. The codebase is at https://github.com/IBM/activated-lora.
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands fine-tuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA) (Hu et al., 2021) has addressed these challenges by introducing a parameter-efficient method for fine-tuning (Houlsby et al., 2019), enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a lightweight and effective alternative for task-specific customization, particularly for large foundation models. Yet, while LoRA excels in static or single-task scenarios, it presents inefficiencies when applied in multiturn interactions, where dynamic switching between multiple LoRA configurations is required in agentic or reasoning pipelines.
@@ 1.1 Background
This lowers the number of parameters, making finetuning significantly more efficient (Hu et al., 2021). If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then the KV cache for the input must be recomputed.
@@ 2 Intrinsics and late prompts
early promptinginstructioncontentanswer
late promptingcontentinstructionanswer
Figure 2: Late vs. early prompting framework for intrinsics. The aLoRA adapter architecture is designed to preserve the cache-reuse benefits of late prompting by adapting weights only on the red tokens, allowing it to reuse the base model cache for the blue input tokens.
@@ 3.1 Invocation
3.2 Training and Huggingface Implementation
Our implementation of aLoRA is at https://github.ibm.com/Kristjan-H-Greenewald/activated-lora/tree/main. It seamlessly supports standard Huggingface training libraries Wolf et al. (2019) as well as inference/generation methods (with or wit hout using available base model KV cache) if needed (e.g. for testing or proofs of concept when the efficiencies of vLLM Zhu et al. (2023) are not needed). Later in this section, we will also show inference experiments using our modification of the more efficient (SOTA) inference package vLLM to support aLoRA.
@@ 3.3 Model Capacity and Increased Rank
3.4 Comparison to LoRA at inference time
In Figure 4 we illustrate how computation and memory differ in LoRAs versus aLoRA in a simple agent pattern. A prompt (blue) is passed to a model which then prefills the corresponding KV cache, and then generates an answer (green). The task is now to evaluate the answer; in this example, two different hypothetical evaluators are used. An example of such evaluations may be to determine if the answer is faithful to the content provided, or how certain the model is of the answer, given the content. If these evaluators are implemented using LoRAs (Figure 4a), then to generate the evaluations, new KV cache computations need to be performed, independently for each evaluator. In the case of aLoRAs (Figure 4b), the KV cache of the underlying model is reused, resulting in significant savings.
To prove this point experimentally, we modified vLLM Zhu et al. (2023) to be able to do inference on activated LoRAs, trained four regular and activated LoRAs, and created a synthetic benchmark where the prompt length and the number of evaluators is varied. In Figure 5 we see the results of the experiment for the setting where the answer is comprised of 256 tokens and each evaluation is given by 16 tokens. The plots depict the total time processing the input prompts divided by the number of such prompts for IBM’s Granite 3.2 8 billion parameter instruct model (Granite 3.2 8b Instruct).
Note that the performance of activated LoRA changes very little as one adds more evaluators (more aLoRA adapted models in parallel), but the performance of regular LoRAs drops significantly as more adapters are added. This trend is consistent across all prompt lengths, showing very significant savings when using activated LoRA in this regime. We also emphasize that the use of aLoRA as an evaluator is for illustrative purposes; aLoRAs can also be used to generate steps in a reasoning process or a final answer in the case of more complex agent applications.
(a)promptanswereval1eval2prefill (base)gen (base)prefill (LoRA1)prefill (LoRA1) gen (LoRA1)prefill (LoRA2)prefill (LoRA2) gen (aLoRA2)(b)promptanswereval1eval2prefill (base)gen (base)gen (aLoRA1)gen (aLoRA2)
Figure 4: Computation and memory pattern of (a) LoRA vs. (b) aLoRA used as evaluators of an answer given by a base model. Narrow rectangles denote tokens and wide rectangles denote the KV cache. Significant GPU compute and memory savings can be realized because aLoRA can reuse the base model cache.
Figure 5: Comparison of aLoRA and LoRA as one varies the number of adapters, for a simple agent pattern. The experiment exemplifies that aLoRA scales much more efficiently than LoRA for this task.
@@ 5 Conclusion
Future work will explore developing expanded use cases for aLoRA models in agentic and test-time-scaling applications and move towards a release of code for vLLM integration.
@@ B.1 QA datasets for Uncertainty Quantification Intrinsic
B.2 Training data for Query Rewrite Intrinsic
@@
B.3 Training data for Answerability Determination Intrinsic